answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Find the probability of rolling a sum of 5 with two dice?

The probability of rolling a sum of 5 with two dice is 4 in 36, or 1 in 9, or about 0.1111.

1 and 4

2 and 3

3 and 2

4 and 1

Find a formula for the probability distribution of the random variable x representing the outcome when a single die is rolled once?

Possible Outcomes when die is rolled

dot showed by die , possible outcome in single roll

  • 1,1
  • 2,1
  • 3,1
  • 4,1
  • 5,1
  • 6,1

thus , formula for the probability distribution of the random variable x will be

P(X=x) = x/6Cx Where as x = 1

What is a space diagram in maths probability?

A space diagram is commonly used in mathematics. It is a

table which represents a range of work to mostly do with probability!

Hope it helps

What is the probability of 2 heterozygous individuals producing a heterozygous recessive offspring?

The probability is 50%.

There are four probabilities: dominant homozygous, recessive homozygous, or heterozygous.

There are 7 things in a hat how many ways can you pick 3 things from the hat at once?

If you are drawing three things out at once, there are 35 combinations.

You solve this with a tree diagram. Consider the items as A,B,C,D,E,F.There are 15 combinations that have A, 10 that have B (not counting the ones with A), 6 with C (but not A or B), 3 with D (but not A,B, or C) and one EFG (without A,B,C or D).

If two coins are tossed what is the possible outcome?

When two coins are tossed, the results that can be achieved are HH, HT, TH, and TT where H represents a heads and T represents a tails. The probability of getting two heads is 1/4, and by symmetry the probability of getting two tails is 1/4. The probability of getting one heads and one tails in some order is 1/2 which is the most likely outcome.

Who is asking these questions anyway? Why cant we insert E for edge in the equation. Would a coin NEVER land on its' edge? How about in the water? Didn't George attempt to throw a dollar across the Delaware River? Did we ever hear if he made it? If it landed in the muddy bank, odds are better it landed on its edge. If someone keeps asking these vague questions, no one can ever give a definitive answer unless a lot of stupid assumptions are made. Now they won't be stupid if you first list them. Now, let's see your list.

A pair of dice is constructed so that each die is marked with a 1 on one side a 2 on two sides and a 3 on three sides The dice are rolled Find the probability that the same number appears on each?

Let X be the number that is rolled

P(X=1) = 1/6

P(X=2) = 2/6

P(X=3) = 3/6

Case 1 (2 ones are rolled)

(P(X=1))(P(X=1)) = 1/36

Case 2 (2 twos are rolled)

(P(X=2))(P(X=2)) = 4/36

Case 3 (2 threes are rolled)

(P(X=3))(P(X=3)) = 9/36

Probability of same number appearing on each =

(Case 1)U(Case 2)U(Case 3) since they are independent this equals

Case1 + Case2 +Case3=

1/36 +4/36 + 9/36 = 14/36

What does it means if the standard deviation is small?

Simplest terms, it means that the sample values are grouped closely together. Another way of saying it is that the range of values is small.

The plot will be a leptokurtic plot (peaked).

What is the probability of a composite number if you rolled a six sided dice?

The probability of rolling a composite number on a 6 sided die, numbered 1, 2, 3, 4, 5, and 6, is 1 in 3.

There are two composite numbers on the die, 4 and 6. The others are prime, 2, 3, and 5, or unity, 1, so there are 2 possible outcomes in an possible result set of 6, hence the probability is 2 in 6, or 1 in 3.

What is the z score?

z = (x - u)/(standard dev)

The z score expresses the difference of the experimental result x from the most probable result u as a number of standard deviations. The probability can then be calculated from the cumulative standard normal distribution. ie sigma(z)

Suppose a population consisted of 20 items How many different samples of n equals 3 are possible?

There are 20 ways to choose the first sample, which leaves 19 ways for the second one, and 18 for the third one. So (apparently) there are 20 x 19 x 18 ways to choose a sample of 3 items, BUT, since the order of choosing a particular set of 3 doesn't matter, we need to divide by 3 x 2 x 1; so the final calculation is (20 x 19 x 18) divided by (3 x 2 x 1). Cancelling by the 3 and the 2 we have 10 x 19 x 6 = 10 x 114 = 1140

Probability of 3 marbles 2 yellow and 4 red in a bag?

The probability depends on the availability of marbles of those colours to the person making up the bag.

What important events occurred during 2009?

See related link. Other people may not view the same events as important.

Would a small deviation in an experiment mean that something is wrong with the experiment?

Not necessarily. It might mean that the experiment has a highly stable outcome. You need to evaluate if that is true or if the experiment is flawed. It comes down to theoretical expectations versus experimental outcomes - you should know a priori (before the fact) what to expect, so you can know if the results are good.

For instance...

If you were measuring the radioactivity of a sample with a relatively low count rate using a detector that recorded counts in each second, you would expect a poissen distribution.

If you were measuring the same sample with a detector that counted for 1 minute, you would expect a more gaussian distribution.

If, on the other hand, you were measuring the wavelength of a red laser, you would expect that every single observation would give you the same results, within an extremely tight distribution.

Trending Questions
How many marbles are in a set? Why is it hard to calculate theoretical probability for bowling? Probability of getting married and having children? How many ways can you put 8 mutually non-attacking rooks on a standard 8x8 chessboard? How often does an biannually event occur? What is the probability of obtaining four tails in five flips of a coin? What is the probability that a diamond will not be pulled out of a deck of cards? In which sport do you have a coin toss? Are mountains discrete or continuous? How many combinations would it take to play every possible outcome to the ncaa bracket? What are the odds in favor of drawing a card higher than a jack from an ordinary deck of cards? The variation in the population as measured by the standard deviation has little or no effect in determining the size of a sample selected from the population True or False? What is the probability of rolling a 4 or an even number on a number cube? How many permutations are there of the letters in the word greet? In what year did in Jim thorpes life in 1929 the biggest event happen? A bag contains 2 red marbles 3 blue marbles and 5 green marbles. if one marble is drawn from the bag what is the chance that the marble willbe red? How much 2.2 crores? Adjust your vehicle to help prevent a dangerous event from occurring or to minimize the consequences of an unavoidable crash.? Division multiplication and factorization in algebra? What is the probability of the die landing on an even number?